OpenStudy (crashonce):

If sinAcosA = 1/3, then 1/sinA^2 + 1/cosA^2 = ? @satellite73

OpenStudy (xapproachesinfinity):

start off by manupilating 1/sin^2+1/cos^2 if you did common denominator you should get (cos^2A+sin^2A)/sinAcosA and we know that cos^2+sin^2=1 and they gave us sinAcosA=1/3

OpenStudy (xapproachesinfinity):

just put things in their place and you got what you are looking for

OpenStudy (xapproachesinfinity):

eh that denominator should be sin^2Acos^2A

OpenStudy (xapproachesinfinity):

hehe i just realize i didn't square them lol

OpenStudy (xapproachesinfinity):

if you are wondering how to get squares just square both sides of sinAcosA=1/3 meaning (sinAcosA)^2=(1/3)^2 which is the same as sin^2Acos^2A=1/9

OpenStudy (xapproachesinfinity):

you have everything now just plug in the values

OpenStudy (crashonce):

wait so its 1/3? @xapproachesinfinity

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\dfrac{1}{sin^2A}+\dfrac{1}{cos^2A}\\~\\ =\dfrac{cos^2A+sin^2A}{sin^2Acos^2A}\\~\\ =\dfrac{1}{(sinAcosA)^2}\\~\\ }\end{align}\)